OpenStudy (anonymous):

James is offered a job that pays $800 for 4 days of work. He is offered another job that pays $950 for 5 days of work. Which job will pay him more per day? the job that pays $800 for 4 days of work the job that pays $950 for 5 days of work Both jobs will pay the same amount per day.

OpenStudy (anonymous):

@ortiz09

OpenStudy (anonymous):

800\4 is 200 950\5 is 190

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

A turtle moved 1 meters in 3 hours. What is the turtle’s speed? \frac{1}{2} \frac{1}{2} \frac{1}{2} \frac{1}{2}

OpenStudy (anonymous):

distance divided by time